# Active nematic pumps

### v1

@@ II.3 Micro-fluidic transport and mixing by active nematic pumps
On the other hand, channel permeability represents an enormous advantage with respect to classical microfluidic devices, and allowed us to achieve enhanced downstream mixing. We tested this concept by simulating an active pump system where adjacent circular ensembles of two distinctive colloidal species (red and blue in Fig. 3f) are advected by active flows. We observe that, while the center of mass of the ensemble moves downstream, mixing of the transported particles takes place, both in the longitudinal direction within channels and between neighboring channels. In fact, most particles move to a different channel during the process, as demonstrated in Fig. Active nematic pumps, where we show that the probability of a particle remaining in its original channel drops steadily with time. The effectiveness of mixing is exemplified in Fig. 3g with the temporal evolution of a mixing parameter, $\chi$, which drops from $\chi=1$ (unmixed) to $\chi\simeq 0.1$ (near perfect mixing) after the ensemble has crossed $\sim 11$ pumping elements. Although mixing is more efficient in fully turbulent active flows performing free of lateral constraints, our pumping system combines mixing with downstream transport, something that cannot be achieved through a lattice of isotropic obstacles or inside a channel with smooth walls (Fig. Active nematic pumps).
@@ II.4 Integrating active pumps into microfluidic systems
The versatility of our active nematic pumping system enables its incorporation anywhere and at anytime within an active fluid layer. As an example, in Fig. 4 we demonstrate the efficacy of our control strategy by directing the flow from the outlet of an AN-pumped virtual channel into a channel confined by physical walls. It is known that flow inside a channel narrower than the active length scale can lead to structured flow patterns and even to net transport, although the resulting flow orientation is randomly selected upon symmetry breaking [23]. In the present case, the orientation of the flow is determined by the active pumping system. Here, both triangular inclusions for pumping and the continuous walls of the channel were fabricated simultaneously. One could, however, incorporate an active pump to feed a preexisting channel. In this system, we observe that the rectified flow speed decays downstream inside the channel, but it maintains the directionality, which is imposed by the active pumps. In this example, the length of the channel is similar to that of the active pump system, and is limited by the field of view in our current experimental setup. We have confirmed the experimental results using numerical simulations with a channel much longer than the size of the AN pumping unit (Fig. 4c). We find that the average velocity in the center of the channel stabilizes to the value at the output from the pump (Fig. 4c). In these simulations, the channel is just slightly wider than the flow lane of the injecting pump system. For wider channels, we observe a decay of the speed, similar to the one observed in experiments (Fig. 4d). Results are different if the output of the active pump system is directed towards an arrangement of non-pumping obstacles, such as an array of circular columns (Fig. Active nematic pumps). The permeable channel walls lead to a quick randomization of the velocity.
Finally, we designed a system that sustained a steady active flow circulation along channels of arbitrary shape. As an example, in Fig. 4f we have simulated a large rectangular loop of active pumps. The average clockwise flow circulation along the center of the channel (rectangular path in the figure) is shown to fluctuate around a steady value for the duration of the simulations (Fig. 4g). Interestingly, the active flows in the region surrounded by the loop of active pumps are no longer isotropic, which provides us with a tool to generate alternative flow patterns. For instance, in Fig. 4h, we show the average flow field created experimentally with a simple ring of elementary active pumps, revealing that the central region of the AN layer has developed a steady counter clockwise flow.
@@ III Discussion
Amidst this complexity, we have demonstrated here a versatile protocol to locally control the directionality of the flows of a microtubule/kinesin AN. By incorporating photopolymerizable reactants in the aqueous medium, we have polymerized lattices of columns with triangular cross section, which force an adaptation of the surrounding active flows, breaking the fore-aft symmetry and transforming the usual isotropic active turbulence regime into ordered arrays of counter-rotating vortices that organize the AN flows along virtual channels, as material flux is allowed across them. Combining experiments and numerical simulations, we have analyzed both the flow velocity and hydrostatic pressure profiles, demonstrating that the new structures behave as classical hydraulic pumps when a small number of elements are created sequentially, with a velocity and hydrostatic pressure that are proportional to the number of pumping elements.
@@ Acknowledgements
We thank M. Pons, A. LeRoux, and G. Iruela for their assistance in the expression of motor proteins. I.V.-C. acknowledges funding from Generalitat de Catalunya though a FI-2020 PhD. Fellowship. P.G. acknowledges support from the European Union’s Horizon Europe Research and Innovation Programme through the Marie Sklodowska-Curie Actions (grant agreement 101065794). I.V.-C., J.I.-M., and F.S. acknowledge funding from MICIU/AEI/10.13039/501100011033 (Grant No. PID2022-137713NB-C21). R. C. and M.T.G. acknowledge financial support from the Portuguese Foundation for Science and Technology (FCT) under the contracts: PTDC/FISMAC/5689/2020 (DOI 10.54499/PTDC/FIS-MAC/5689/2020), EXPL/FIS-MAC/0406/2021 (DOI 10.54499/EXPL/FIS-MAC/0406/2021), UIDB/00618/2020 (DOI 10.54499/UIDB/00618/2020), UIDP/00618/2020 (DOI 10.54499/UIDP/00618/2020), DL 57/2016/CP1479/CT0057 (DOI 10.54499/DL57/2016/CP1479/CT0057) and 2023.10412.CPCA.A2. The authors are indebted to the Brandeis University MRSEC Biosynthesis facility, supported by NSF MRSEC 2011846, for providing the tubulin.
@@ III.1.3 Imaging and photopatterning
The fluorescence microscope was modified in order to incorporate a Ti Light Crafter 4500 DLP development modules (EKB Technologies, Ltd.) equipped with a high 2 W 385 nm LED source in one case. Projected patterns are incorporated into the light path of the inverted microscope by means of a collimating lens (f = +150 mm) and a 505 nm dichroic mirror (Thorlabs DMLP505R), and are focused on the sample by means of the microscope objectives, reaching a lateral resolution up to a few microns. The DMD projector is connected as an external monitor to a computer, thus enabling real time control of the projected patterns using MS-PowerPoint slides. Hydrogel objects were polymerized using the x10 objective, resulting in a light power density of 1.6 W cm-2), and a irradiation time of 2 seconds.
III.1.4 PIV analysis
@@
The flow field of the active material was extracted by means of particle image velocimetry (PIV) from the fluorescent images of the Dragon Green particles dispersed in the material using the public MatLab App PIVLab.
III.2 Simulations
@@
We model the active nematics using a 2D continuum hydrodynamic theory which is summarized in this section. The director field $\mathbf{n}(\mathbf{x})$ gives the average orientation of the filaments in the AN at the position $\mathbf{x}$ while the degree of order is given by the scalar order parameter $S$, which is $S=0$ in the isotropic phase and has a finite value $S_{N}$ in the nematic phase. Because the filaments have a head-tail symmetry, a tensor order parameter is used: $Q_{\alpha\beta}=2S\left(n_{\alpha}n_{\beta}-\delta_{\alpha\beta}/2\right)$. In the free energy density, the bulk terms are not considered. It is assumed that the fluid becomes nematic only when it is active, which mimics the behaviour of the mixtures of microtubule and kinesin [37]. The free energy density includes an elastic term responsible for penalizing distortions in the order parameter, $f_{E}=K\left(\partial_{\gamma}Q_{\alpha\beta}\right)^{2}/2$, where $K$ is the elastic constant, and an anchoring term applied only at the solid surfaces, $f_{W}=W(Q_{\alpha\beta}-Q_{\alpha\beta}^{0})^{2}/2$, which penalises deviations from the surface-preferred order parameter $Q_{\alpha\beta}^{0}$ with strength $W$. The total free energy is the sum of these two terms integrated in the entire area: $\mathcal{F}=\int fd^{2}x$, where $f=f_{E}+f_{W}$ at the solid surface and $f=f_{E}$ elsewhere. The dynamics is governed by the Beris-Edwards (for the tensor order parameter $Q_{\alpha\beta}$), continuity and Navier-Stokes equations (for the velocity field $\mathbf{u}$), respectively [38, 39, 20]:
@@
In Eq. (1), $\Gamma$ stands for the rotational diffusivity, $H_{\alpha\beta}$ for the molecular field: $H_{\alpha\beta}=-\frac{\delta\mathcal{F}}{\delta Q_{\alpha\beta}}+\frac{\delta_{\alpha\beta}}{2}\operatorname{Tr}\left(\frac{\delta\mathcal{F}}{\delta Q_{\gamma\epsilon}}\right)$ and $S_{\alpha\beta}$ for the co-rotational term: $S_{\alpha\beta}=(\xi D_{\alpha\gamma}+W_{\alpha\gamma})\left(Q_{\beta\gamma}+\delta_{\beta\gamma}/2\right)+\left(Q_{\alpha\gamma}+\delta_{\alpha\gamma}/2\right)(\xi D_{\gamma\beta}-W_{\gamma\beta})-2\xi\left(Q_{\alpha\beta}+\delta_{\alpha\beta}/2\right)(Q_{\gamma\epsilon}\partial_{\gamma}u_{\epsilon})$, where the vorticity and shear rate are $W_{\alpha\beta}=(\partial_{\beta}u_{\alpha}-\partial_{\alpha}u_{\beta})/2$ and $D_{\alpha\beta}=(\partial_{\beta}u_{\alpha}+\partial_{\alpha}u_{\beta})/2$. The aligning parameter $\xi$ controls the relative importance of the vorticity and shear rate and it is related to the shape of the particles that compose the nematics: rod-like particles have $\xi>0$, while disk-like particles have $\xi<0$. We consider rod-like particles (filaments) in the flow-aligning regime. In Eq. (3), $\rho$ is the fluid density, $p$ is the fluid pressure and $\eta$ is the absolute viscosity. The activity level is controlled by the parameter $\zeta$, which is positive for extensile systems as those in this work. The equations described above are a simplified version of a standard 2D model for active nematics [37, 40, 20]. Because the active stresses are typically much larger than the passive ones [29], we neglected the passive stress in Eq. (3) as in previous works [22, 41].
@@
A hybrid numerical method is used in the simulations. Equation (1) is solved using a predictor-corrector finite-differences method while Eqs. (2) and (3) are recovered in the macroscopic limit using the lattice Boltzmann method [42, 40]. At the solid boundaries, we use planar weak anchoring and no-slip boundary conditions by means of half-way bounce-back boundary conditions. The initial conditions are the fluid at rest and directors mostly aligned with the horizontal direction with random fluctuations of ±$\pm 2^{\circ}$. Different seeds are used to generate random initial conditions. The parameters used in the simulations are as follows (except where otherwise stated) in lattice units, in which the reference density, lattice spacing and time step are equal to one: absolute viscosity $\eta=6.67$, elastic constant $K=0.015$, aligning parameter $\xi=0.9$, activity $\zeta=0.01$, fluid density $\rho=40$, anchoring strength $W=0.002$, rotational diffusivity $\Gamma=0.4$, obstacle size $H=70$ and size of the square unit cell $L=128$.

### v2

@@ II.3 Micro-fluidic transport and mixing by active nematic pumps
On the other hand, channel permeability represents an enormous advantage with respect to classical microfluidic devices, and allowed us to achieve enhanced downstream mixing. We tested this concept by simulating an active pump system where adjacent circular ensembles of two distinctive colloidal species (red and blue in Fig. 4f) are advected by active flows. We observe that, while the center of mass of the ensemble moves downstream, mixing of the transported particles takes place, both in the longitudinal direction within channels and between neighboring channels. In fact, most particles move to a different channel during the process, as demonstrated in Fig. Active nematic pumps, where we show that the probability of a particle remaining in its original channel drops exponentialy with time. The effectiveness of mixing is exemplified in Fig. 4g with the temporal evolution of a mixing parameter, $\chi=\langle(N_{A}-N_{B})^{2}\rangle/\text{max}\langle(N_{A}-N_{B})^{2}\rangle$, where $N_{A}$ and $N_{B}$ are the number of particles of the two species at each position (within a square of side half the cell size). Perfect mixing corresponds to $\chi=0$. We observe a drop from $\chi=1$ (unmixed) to $\chi\simeq 0.1$ (near perfect mixing) after the ensemble has crossed $\sim 11$ pumping elements. Although mixing is more efficient in fully turbulent active flows performing free of lateral constraints, our pumping system combines mixing with downstream transport, something that cannot be achieved through a lattice of isotropic obstacles or inside a channel with smooth walls (Fig. Active nematic pumps). To put the performance of our active pumps in the context of microfluidic mixing [38], we have computed a Péclet number, $Pe=vL/D$, which compares particle transport by advection and by diffusion [39]. To this end, we have extracted an effective diffusion coefficient along the transverse direction and used, as advection velocity, the average speed along the virtual channels (Fig. Active nematic pumps). In experiments, we have $\langle|v_{x}|\rangle\simeq 1\,\mu$m/s, $L\simeq 100\,\mu$m, and $D_{eff}=15.2\,\mu m^{2}/s$, which yields $Pe\simeq 6.6$. In simulations, we have $\langle|v_{x}|\rangle\simeq 0.0042\,l.u.$, $L\simeq 128\,l.u.$, and $D_{eff}=0.077$ l.u., which gives $Pe\simeq 7.0$, consistent with experiments. $Pe$ values close to unity, such as the ones obtained here, indicate that diffusion and advection times are comparable for the transport of dispersed particles. Note that $D_{eff}$ is several orders of magnitude larger than $D$, the thermal diffusion coefficient, as the bulk viscosity of the AN layer is $\eta\simeq 1$ Pa·s [29]. The observed diffusion enhancement is an example of active Taylor dispersion [39], a phenomenon amply studied in passive liquids, but that has been addressed only recently for active fluids [40, 41].
@@ II.4 Integrating active pumps into microfluidic systems
The versatility of our active nematic pumping system enables its incorporation anywhere and at anytime within an active fluid layer. As an example, in Fig. 5a we demonstrate the efficacy of our control strategy by directing the flow from the outlet of an AN-pumped virtual channel into a channel confined by physical walls. It is known that flow inside a channel narrower than the active length scale can lead to structured flow patterns and even to net transport, although the resulting flow orientation is randomly selected upon symmetry breaking [23]. In the present case, the orientation of the flow is determined by the active pumping system. Here, both triangular obstacles for pumping and the continuous walls of the channel were fabricated in-situ simultaneously. One could, however, incorporate an active pump to feed a preexisting channel. In this system, we observe that the rectified flow speed decays downstream inside the channel, but it maintains the directionality, which is imposed by the active pumps (Fig. 5b). In this example, the length of the channel is similar to that of the active pump system, and is limited by the field of view in our current experimental setup. We have confirmed the experimental results using numerical simulations with a channel much longer than the size of the AN pumping unit (Fig. 5c). We find that the average velocity in the center of the channel stabilizes to the value at the output from the pump (Fig. 5c). In these simulations, the channel is just slightly wider than the flow lane of the injecting pump system. For wider channels, we observe a decay of the speed, similar to the one observed in experiments (Fig. 5d). The effect of channel width on the sustained horizontal velocity after injection is further explored using simulations in Fig. Active nematic pumps. For the same channels, and in the absence of active injection, the average horizontal velocity fluctuates around zero.
The pumped flows can be sustained within long channels, as demonstrated with larger scale simulations in Fig. Active nematic pumps. Results are different if the output of the active pump system is directed towards an arrangement of non-pumping obstacles, such as an array of circular columns (Fig. Active nematic pumps), where the permeable channel walls lead to a quick randomization of the velocity.
Our strategy allows to impose a steady active flow circulation along channels of arbitrary shape. As an example, in Fig. 5e for experiments, and in Fig. 5f for the corresponding simulations, we show active flows generated by a large square loop of active pump elements. Interestingly, the active flows in the region surrounded by the loop of active pumps are no longer isotropic, which provides us with a tool to generate alternative flow patterns normally not found in active nematics. In this example, the average counterclockwise flow field created by the ring of active pumps develops a steady counterclockwise flow in the central region surrounded by triangular obstacles. We have monitored the evolution of the flow circulation along the center of the square channel (Fig. Active nematic pumps), and have observed that a steady state is established after a brief transient. Besides the versatility demonstrated by these examples, our approach offers real time reconfigurability of the fluidic devices. In Movie S7 (also simulations in Movie S8) we present experiments where a simple array of active pumps is imprinted, organizing a rightwards average flow inside the virtual channels. Then, we imprint larger triangular obstacles on top of the original ones, this time rotated 180 degrees. After a short transient, the flow is reversed by the new array of pumps. This can be observed in Fig. 5g and 5h, where the evolution of the average horizontal velocity is represented for experiments and simulations, respectively.
@@ III Discussion
Amidst this complexity, we have demonstrated here a versatile protocol to locally control the directionality of the flows of a microtubule/kinesin AN. By incorporating photopolymerizable reactants in the aqueous medium, we have polymerized lattices of columns with triangular cross section, which force an adaptation of the surrounding active flows, breaking the fore-aft symmetry and transforming the usual isotropic active turbulence regime into ordered arrays of counter-rotating vortices that organize the AN flows along virtual channels, as material flux is allowed across them. Combining experiments and numerical simulations, we have analyzed both the orientational field, the flow velocity, and hydrostatic pressure profiles. We have identified a pattern in the time-averaged defect arrangement that demonstrates that the physical mechanism at the origin of active pumping is the modification of the defect dynamics due to the interaction of the active nematic with the inclusions. Our study has also demonstrated that the new structures behave as classical hydraulic pumps when a small number of elements are created sequentially, with a velocity and hydrostatic pressure that are proportional to the number of pumping elements.
@@ Acknowledgements
We thank M. Pons, A. LeRoux, and G. Iruela for their assistance in the expression of motor proteins. I.V.-C. acknowledges funding from Generalitat de Catalunya though a FI-2020 PhD. Fellowship. P.G. acknowledges support from the European Union’s Horizon Europe Research and Innovation Programme through the Marie Sklodowska-Curie Actions (grant agreement 101065794). I.V.-C., J.I.-M., and F.S. acknowledge funding from MICIU/AEI/10.13039/501100011033 (Grant No. PID2022-137713NB-C21). I.V.-C. and J.I.-M. acknowledge funding from MICIU/AEI/10.13039/501100011033 (Grant No. PDC2022-133625-I00). R. C. and M.T.G. acknowledge financial support from the Portuguese Foundation for Science and Technology (FCT) under the contracts: PTDC/FISMAC/5689/2020 (DOI 10.54499/PTDC/FIS-MAC/5689/2020), EXPL/FIS-MAC/0406/2021 (DOI 10.54499/EXPL/FIS-MAC/0406/2021), UIDB/00618/2020 (DOI 10.54499/UIDB/00618/2020), UIDP/00618/2020 (DOI 10.54499/UIDP/00618/2020), DL 57/2016/CP1479/CT0057 (DOI 10.54499/DL57/2016/CP1479/CT0057) and 2023.10412.CPCA.A2. The authors are indebted to the Brandeis University MRSEC Biosynthesis facility, supported by NSF MRSEC 2011846, for providing the tubulin.
Author contribution. I.V.-C. performed most experiments and data analysis, with collaboration from P. G.; M. V. performed and analyzed some of the experiments; R.C.V.C. performed the numerical simulations, with collaboration from M.T.G.; J. I.-M. and F. S. supervised the research; J. I.-M. wrote the manuscript with collaboration from all authors.
Data availability. All study data are included in the article and/or supporting information. The data can also be accessed at: doi: 10.17605/OSF.IO/N8G4C. [43]
Competing interests. The authors declare no competing interests.
@@ III.1.3 Imaging and photopatterning
The fluorescence microscope was modified in order to incorporate a Ti Light Crafter 4500 DLP development modules (EKB Technologies, Ltd.) equipped with a high 2 W 385 nm LED source in one case. Projected patterns are incorporated into the light path of the inverted microscope by means of a collimating lens (f = +150 mm) and a 505 nm dichroic mirror (Thorlabs DMLP505R), and are focused on the sample by means of the microscope objectives, reaching a lateral resolution up to a few microns. The DMD projector is connected as an external monitor to a computer, thus enabling real time control of the projected patterns using MS-PowerPoint slides. Hydrogel objects were polymerized using the Nikon CFI Plan Fluor DLL x10 objective (N.A. = 0.3), resulting in a light power density of 1.6 W cm-2), and a irradiation time of 2 seconds. The imprinted patterns are currently confined within the field of view, which is around $850\,\mu$m×$\times 650\,\mu$m in the used optical setup. The studied active pump systems are in all cases surrounded by extended regions, at least 1 mm in all directions, where the unconstrained active nematic is in the active turbulent regime.
III.1.4 Flow field analysis and particle tracking
@@
The flow field of the active material was extracted by performing particle image velocimetry directly on the fluorescence images using the public domain MatLab App PIVLab. To study the transport of advected colloidal particles, we have included a small amount of $1\,\mu$m hydrophilic fluorescent polystyrene microparticles (Dragon Green, FCDG006, Bangs Labs), and tracked their motion from fluorescence images using the plugin Particle Tracker 2D/3D in FIJI.
III.2 Orientational field analysis and defect detection
The local active nematic tensor field was obtained using a coherence-enhanced diffusion filtering (CEDF) MatLab code [44] applied on fluorescence images of the AN layer. This analysis yields the local scalar order parameter and director field, and allows to detect the topological defects. The local topological charge density, $\sigma$, is obtained from the director field, $n_{i}$ by defining the tensor $q_{ij}=n_{i}n_{j}-1/2\delta_{ij}$ and $\sigma=-\frac{1}{\pi}\left(\frac{\partial q_{11}}{\partial x}\frac{\partial q_{12}}{\partial y}+\frac{\partial q_{12}}{\partial x}\frac{\partial q_{22}}{\partial y}-\frac{\partial q_{12}}{\partial x}\frac{\partial q_{11}}{\partial y}-\frac{\partial q_{22}}{\partial x}\frac{\partial q_{12}}{\partial y}\right)$.
III.3 Simulations
@@
We model the active nematics using a 2D continuum hydrodynamic theory which is summarized in this section. The director field $\mathbf{n}(\mathbf{x})$ gives the average orientation of the filaments in the AN at the position $\mathbf{x}$ while the degree of order is given by the scalar order parameter $S$, which is $S=0$ in the isotropic phase and has a finite value $S_{N}$ in the nematic phase. Because the filaments have a head-tail symmetry, a tensor order parameter is used: $Q_{\alpha\beta}=2S\left(n_{\alpha}n_{\beta}-\delta_{\alpha\beta}/2\right)$. In the free energy density, the bulk ordering terms are not considered. This corresponds to placing the passive nematic at the instability of the isotropic phase, the isotropic spinodal. At this point the passive system is unstable to nematic fluctuations. Our motivation for this choice of parameters arises from experimental observations in microtubule-kinesin mixtures, where the nematic order emerges driven by activity alone. While including bulk terms modifies the stability of the passive system, these terms are not expected to affect the dynamical steady state qualitatively, as activity drives both the passive ordered state and the disordered state considered here linearly unstable with a zero threshold for large systems. A similar approach has been considered in related studies [45, 22, 46]. The free energy density includes an elastic term responsible for penalizing distortions in the order parameter, $f_{E}=K\left(\partial_{\gamma}Q_{\alpha\beta}\right)^{2}/2$, where $K$ is the elastic constant, and an anchoring term applied only at the solid surfaces, $f_{W}=W(Q_{\alpha\beta}-Q_{\alpha\beta}^{0})^{2}/2$, which penalises deviations from the surface-preferred order parameter $Q_{\alpha\beta}^{0}$ with strength $W$. The total free energy is the sum of these two terms integrated in the entire area: $\mathcal{F}=\int fd^{2}x$, where $f=f_{E}+f_{W}$ at the solid surface and $f=f_{E}$ elsewhere. The dynamics is governed by the Beris-Edwards (for the tensor order parameter $Q_{\alpha\beta}$), continuity and Navier-Stokes equations (for the velocity field $\mathbf{u}$), respectively [47, 48, 20]:
@@
In Eq. (1), $\Gamma$ stands for the rotational diffusivity, $H_{\alpha\beta}$ for the molecular field: $H_{\alpha\beta}=-\frac{\delta\mathcal{F}}{\delta Q_{\alpha\beta}}+\frac{\delta_{\alpha\beta}}{2}\operatorname{Tr}\left(\frac{\delta\mathcal{F}}{\delta Q_{\gamma\epsilon}}\right)$ and $S_{\alpha\beta}$ for the co-rotational term: $S_{\alpha\beta}=(\xi D_{\alpha\gamma}+W_{\alpha\gamma})\left(Q_{\beta\gamma}+\delta_{\beta\gamma}/2\right)+\left(Q_{\alpha\gamma}+\delta_{\alpha\gamma}/2\right)(\xi D_{\gamma\beta}-W_{\gamma\beta})-2\xi\left(Q_{\alpha\beta}+\delta_{\alpha\beta}/2\right)(Q_{\gamma\epsilon}\partial_{\gamma}u_{\epsilon})$, where the vorticity and shear rate are $W_{\alpha\beta}=(\partial_{\beta}u_{\alpha}-\partial_{\alpha}u_{\beta})/2$ and $D_{\alpha\beta}=(\partial_{\beta}u_{\alpha}+\partial_{\alpha}u_{\beta})/2$. The aligning parameter $\xi$ controls the relative importance of the vorticity and shear rate and it is related to the shape of the particles that compose the nematics: rod-like particles have $\xi>0$, while disk-like particles have $\xi<0$. We consider rod-like particles (filaments) in the flow-aligning regime. In Eq. (3), $\rho$ is the fluid density, $p$ is the fluid pressure and $\eta$ is the absolute viscosity. The activity level is controlled by the parameter $\zeta$, which is positive for extensile systems as those in this work. The equations described above are a simplified version of a standard 2D model for active nematics [45, 49, 20]. Because the active stresses are typically much larger than the passive ones [29], we neglected the passive stress in Eq. (3) as in previous works [22, 50]. While the director field can point out of the plane in 3D models under certain conditions [51], this behaviour is not observed in the present experiments. Therefore, the 2D model appropriately describes the reported phenomena.
A hybrid numerical method is used in the simulations. Equation (1) is solved using a predictor-corrector finite-differences method while Eqs. (2) and (3) are recovered in the macroscopic limit using the lattice Boltzmann method [52, 49]. At the solid obstacle surfaces, we use planar weak anchoring and no-slip boundary conditions in line with known experimental conditions [23], by means of half-way bounce-back boundary conditions. We have verified that, even at zero anchoring strength, the directors align mostly parallel to the obstacle walls. Incrementing the anchoring strength only decreases the average speed. Similarly, we have tested numerical simulations with free-slip boundary conditions of the velocity field on the obstacle walls, finding the same qualitative results. The reported pumping phenomenon, therefore, is robust with respect to the interactions between the active nematic material and the obstacle walls. Simulations are either performed within a closed box, or with periodic boundary conditions in all directions. For closed boxes, the size of the simulation box is adapted to the obstacle system under study, allowing a sufficiently wide surrounding layer of turbulent active nematic. We have verified that this choice of boundary conditions has no relevant impact on the qualitative performance or in the velocity or pressure profiles within the studied pumping systems.
@@
The initial conditions are the fluid at rest and directors mostly aligned with the horizontal direction with random fluctuations of ±$\pm 2^{\circ}$. Different seeds are used to generate random initial conditions. The parameters used in the simulations are as follows (except where otherwise stated) in lattice units, in which the reference density, lattice spacing and time step are equal to one: absolute viscosity $\eta=6.67$, elastic constant $K=0.015$, aligning parameter $\xi=0.9$, activity $\zeta=0.01$, fluid density $\rho=40$, anchoring strength $W=0.002$, rotational diffusivity $\Gamma=0.4$, obstacle size $H=70$ and size of the square unit cell $L=128$.
@@
The tracers are simulated as massless particles in the overdamped regime that follow the fluid velocity and do not collide among themselves. Thus, the tracer’s position is calculated as $\mathbf{x}_{i}(t+\Delta t)=\mathbf{x}_{i}(t)+\mathbf{u}(\mathbf{x})\Delta t$, where $i$ is the index that identifies each tracer and $\Delta t$ is the times step from the lattice Boltzmann simulation.